### v1

# Implementation and Performance Evaluation of TCP over QUIC Tunnels

Abstract
QUIC, a UDP-based transport protocol, addresses several limitations of TCP by offering built-in encryption, stream multiplexing, and improved loss recovery. To extend these benefits to legacy TCP-based applications, this paper explores the implementation and evaluation of a TCP over QUIC tunneling approach. A lightweight, stream-based tunnel is constructed using the Rust-based Quinn library, enabling TCP traffic to traverse QUIC connections transparently. Performance is evaluated under varying network conditions, including packet loss, high latency, and out-of-order delivery. Results indicate that TCP over QUIC maintains significantly higher throughput than native TCP in lossy or unstable environments, with up to a high improvement under 20% packet loss. However, under ideal network conditions, tunneling introduces modest overhead due to encryption and user-space processing. These findings provide insights into the trade-offs of TCP over QUIC tunneling and its suitability for deployment in dynamic or impaired networks.
keywords:
QUIC Protocol, TCP Encapsulation, Transport Layer Tunneling, Network Resilience
1 Introduction
The evolution of transport protocols has been a driving force in the advancement of modern web applications. Traditional protocols like the Transmission Control Protocol (TCP), as defined in RFC 9293 [1] and RFC 793 [2], have long been the backbone of reliable data transmission over the Internet. However, with the increasing demand for low-latency, high-throughput, and secure communication, TCP’s limitations have become more pronounced [3]. TCP’s reliance on a three-way handshake for connection establishment, susceptibility to head-of-line (HOL) blocking, and separate TLS implementation post-handshake introduce inefficiencies that hinder its performance in dynamic network environments.
To address these challenges, the QUIC (Quick UDP Internet Connections) protocol, initially developed by Google and standardized in RFC 9000 [4], has emerged as a promising alternative. QUIC leverages UDP as its transport layer, enabling a 0-RTT handshake for known servers, integrating TLS 1.3 for enhanced security [5], and supporting stream multiplexing to mitigate HOL blocking. These features make QUIC particularly suitable for modern web applications such as HTTP/3, where low latency and high security are critical.
Despite QUIC’s advantages, its adoption in traditional TCP-based applications remains limited [6]. Many legacy systems and applications rely heavily on TCP’s ordered delivery and reliability guarantees. This dichotomy raises the question: how can we harness QUIC’s benefits while maintaining compatibility with TCP-based applications? One potential solution is to encapsulate TCP traffic within QUIC tunnels, leveraging QUIC’s features while preserving the familiar semantics of TCP.
This paper explores the implementation and performance evaluation of a TCP over QUIC tunneling scheme. By encapsulating TCP traffic within QUIC packets, we aim to bridge the gap between traditional TCP applications and modern QUIC-based infrastructure. The study focuses on key performance metrics such as latency, throughput, and reliability across diverse network conditions. Through experimental evaluation, we demonstrate QUIC’s superiority in high-loss and fluctuating network environments, while also addressing challenges such as compatibility issues, computational overhead, and security concerns.
The remainder of this paper is structured as follows. Section 2 provides a detailed comparison of QUIC and TCP core mechanisms, along with an overview of network tunneling principles. Section 3 reviews related work on QUIC and TCP performance studies, as well as network tunneling research. Section 4 describes the implementation details of the TCP over QUIC tunneling scheme. Section 5 presents the experimental methodology. Section 6 presents evaluation metrics, then discusses the results of the performance evaluation. Finally, Section 7 concludes the paper and outlines future research directions.
In this work, we implement a simple TCP over QUIC tunneling system and evaluate its performance under diverse network conditions. Our contributions include a lightweight Rust-based implementation, a thorough comparison with native TCP, and a detailed analysis of throughput, latency, and robustness. The results provide new insights into the feasibility and advantages of using QUIC as a transport layer for TCP-based applications.
2 Background & Motivation
2.1 Comparison of QUIC and TCP core mechanisms
The Transmission Control Protocol (TCP), as defined in RFC 9293 [1] and RFC 793 [2], has been fundamental for reliable Internet data transmission. However, QUIC, initially developed by Google and standardized in RFC 9000 [4], addresses TCP’s limitations, particularly for web applications requiring low latency and high security. A detailed comparison reveals:
  • Connection Establishment: TCP employs a three-way handshake (SYN, SYN-ACK, ACK), requiring at least one RTT, with additional RTT for TLS 1.3 negotiation, as noted in Performance evaluation in TCP and QUIC [7]. QUIC uses UDP as its transport layer, which enables a 0-RTT handshake for known servers. By merging the TCP and TLS handshakes into a single three-packet exchange, QUIC reduces latency and enhances initial data transmission speed.
  • Multiplexing and Streams: TCP operates as a single ordered byte stream, susceptible to head-of-line (HOL) blocking, where packet loss delays all subsequent data. QUIC, as detailed in RFC 9000, supports arbitrary concurrent streams within a connection, each with independent flow control, mitigating HOL blocking and improving resource utilization [8].
  • Security: TCP relies on separate TLS implementation post-handshake (RFC 8446), adding overhead. QUIC integrates TLS 1.3, encrypting all packets by default, reducing handshake latency and enhancing privacy, as discussed in QUIC vs TCP computational efficiency [9].
  • Congestion Control: Both protocols implement congestion control, with TCP using algorithms like New Reno (RFC 6582) [10] and QUIC adopting similar mechanisms (RFC 9002) [11]. However, QUIC’s user-space implementation allows for more agile updates, potentially offering better adaptability to network conditions.
  • Transport Layer Implementation: TCP benefits from kernel-level optimizations, reusing state for efficiency. QUIC, running over UDP in user mode, sacrifices some kernel efficiency for flexibility, as highlighted in performance evaluations showing higher per-packet overhead due to user-kernel boundary crossings [9].
2.2 Network tunneling principle
Network tunneling, involves encapsulating one protocol within another to transport data across incompatible networks or enhance security [12], commonly used in virtual private networks (VPNs). For TCP over QUIC tunnels, the process entails wrapping TCP traffic within QUIC packets, leveraging QUIC’s features like reduced latency and integrated security.
Two modes are relevant [13]:
  • Datagram Mode: This mode, as proposed in Tunneling TCP inside QUIC [14], encapsulates entire packets, including network and transport headers, within QUIC datagrams. It supports any network-layer protocol but incurs significant overhead, with each packet requiring additional IP/UDP/QUIC headers, making it inefficient for TCP, especially with small packets.
  • Stream Mode: Detailed in Tunneling TCP inside QUIC [14], this mode transports TCP bytestreams over QUIC streams, reducing overhead by eliminating redundant headers. It uses Type-Length-Value (TLV) messages for negotiation, such as TCP Connect TLV (0x00, 20 bytes) for destination specification, and supports flow control tuned to TCP’s receive window, enhancing efficiency for applications with frequent small packets. For this reason, we will adopt the stream mode in the implementation.
3 Related Work
3.1 QUIC and TCP Performance Studies
The comparative performance of QUIC and TCP has been extensively studied in recent years, particularly in the context of web and mobile applications requiring low latency, high throughput, and strong reliability. QUIC, as a user-space transport protocol built atop UDP, introduces innovations such as multiplexed streams, integrated TLS 1.3, and 0-RTT connection establishment, positioning it as a successor to TCP for many latency-sensitive applications.
Several studies have also focused on the efficiency of QUIC implementations in practice. They compared the CPU utilization of QUIC and TCP during high-load transmission, showing that QUIC consumes more computational resources due to user-space processing and encryption overhead. Nevertheless, this trade-off is often justified by QUIC’s resilience in lossy or fluctuating networks, where TCP performance degrades rapidly.
In the context of tunneling, encapsulating TCP traffic over QUIC offers a hybrid solution that combines TCP’s widespread application compatibility with QUIC’s transport-layer robustness. This has motivated emerging research into TCP-over-QUIC systems that seek to leverage QUIC’s loss recovery, multiplexing, and security capabilities without altering the semantics of existing TCP applications. Our study extends this line of research by implementing and experimentally evaluating a practical TCP-over-QUIC tunnel using the Quinn library in Rust, further validating the protocol’s strengths and trade-offs under controlled packet loss and latency conditions.
3.2 Comparison of different QUIC implementations
Comparison of different QUIC implementations reveals notable differences in performance and characteristics. According to the work by Salim Hertelli et al. [15], various QUIC implementations such as Chromium-quic ”Quiche”, Cloudflare ”Quiche”, Mvfst, Msquic, Quic-go, Aioquic, and Haskell-quic have been developed in diverse programming languages like C/C++, Rust, Go, Python, and Haskell. These implementations support different versions of QUIC and employ various congestion control algorithms, including CUBIC, NewReno, and HyStart++. While all adhere to IETF specifications, they exhibit significant variations in project status, maintenance, supported features, and performance metrics.
In a more detailed benchmarking study by Kevin Marti [16], MsQuic and Quinn were compared under different networking conditions. The results indicated that increasing the segment size for Quinn can significantly enhance its performance, up to doubling it in some cases. This suggests that Quinn benefits from larger segment sizes, which can reduce the overhead associated with frequent system calls and data copying. In fact, the patched version of Quinn, which sends larger segments, demonstrated improved throughput and fewer system calls compared to its unpatched version. This highlights a potential advantage of Quinn in scenarios where segment size can be optimized. However, when packet loss or reordering was introduced, MsQuic consistently outperformed Quinn by at least 50%. This indicates that MsQuic has more robust mechanisms for handling network instability. Despite this, Quinn’s ability to achieve high performance with optimized segment sizes and reduced system calls suggests that it can be highly efficient under stable network conditions. Overall, while MsQuic demonstrates superior resilience to network instability, Quinn’s performance can be significantly enhanced through specific optimizations, highlighting the importance of implementation-specific tuning in QUIC deployments. Therefore, we will use Quinn’s QUIC implementation to complete the TCP over QUIC tunnel construction.
3.3 Identified Deficiencies
Currently, there is no mature or standardized implementation of TCP over QUIC written in Rust. While some experimental projects exist, they are often incomplete or tailored to narrow use cases. In this work, rather than aiming for a production-ready or fully optimized solution, we present a minimal and functional implementation intended primarily for empirical evaluation. The goal is to provide a foundation for systematic performance testing and to better understand the behavior and trade-offs of TCP over QUIC tunneling under various network conditions.
Moreover, there is a noticeable lack of comprehensive and methodologically sound performance evaluations for TCP over QUIC tunnel architectures. Prior studies have largely focused on comparing QUIC and TCP as standalone transport protocols, without exploring the unique characteristics and trade-offs introduced by encapsulating TCP over QUIC. In particular, metrics such as CPU overhead, throughput under varying packet loss rates, and performance in high-latency environments have not been systematically addressed in the context of tunneling.
This paper aims to fill these gaps by implementing a practical TCP over QUIC tunnel in Rust and conducting targeted performance evaluations across a range of network conditions.
4 System Design & Implementation
Figure 1: Improved Principle of QUIC-Tun: Showing the flow of network traffic from a local TCP client through a QUIC tunnel to a remote TCP server.
This section delineates the implementation details of the TCP over QUIC tunneling. The Figure 1 shows the data flow through the tunnel. The implementation comprises three core components:
  • A data transfer utility facilitating bidirectional communication
  • A server-side QUIC tunnel endpoint
  • A client-side TCP to QUIC proxy
The system was implemented in Rust leveraging the Tokio [17] asynchronous runtime and Quinn [18] QUIC library.
4.1 Bidirectional Data Transfer Mechanism
The core data transfer functionality is encapsulated in the bidirectional_copy utility function. This asynchronous function establishes a full-duplex communication channel between a TCP stream and a QUIC connection by:
  1. 1.
    Splitting the TCP stream into separate reader and writer using tokio::net::TcpStream::split()
  2. 2.
    Creating two parallel asynchronous tasks
    • Continuously copies data from TCP to QUIC send stream until EOF
    • Relays incoming QUIC data to the TCP writer
  3. 3.
    Employing tokio::try_join! macro to synchronize both tasks while propagating any errors
  4. 4.
    Implementing proper stream shutdown procedures upon completion or failure
The implementation ensures atomic operation through Rust’s ownership system and automatic resource management.
4.2 Server-Side
The QUIC server implementation operates as a persistent tunnel endpoint with the following architectural characteristics:
4.2.1 Connection Handling
  • Utilizes Quinn’s Endpoint abstraction to listen for incoming QUIC connections
  • Spawns independent Tokio tasks for each connection to ensure isolation
  • Implements continuous stream acceptance through accept_bi() for bidirectional communication
4.2.2 Configuration Parameters
  • Generates self-signed certificates for development purposes via rcgen.
  • Configures transport parameters including:
    • Maximum 100 concurrent bidirectional streams
    • Disabled unidirectional streams
    • 2-second keep-alive interval
  • Maintains compatibility with standard TLS certificate verification.
4.2.3 Tunnel Establishment
  • For each accepted QUIC stream, establishes a new TCP connection to the target service
  • Delegate data transfer to the bidirectional_copy utility
  • Implements robust error handling for connection timeouts and application-level closures
4.3 Client-Side
The client component functions as a TCP to QUIC proxy with the following implementation details:
4.3.1 Connection Initiation
  • Binds to specified local TCP port using tokio::net::TcpListener
  • Establish QUIC connection to remote endpoint up-on TCP client connection
  • Opens new bidirectional QUIC stream for each TCP connection
4.3.2 Security Configuration
  • Supports configurable certificate verification:
    • Standard TLS certificate chain validation
    • Insecure mode with disabled server verification (for testing)
  • Maintains identical keep-alive parameters (2-second interval) with server
4.3.3 Data Forwarding
  • Instantiates separate Tokio task per connection
  • Utilizes same bidirectional_copy mechanism for consistent data transfer behavior
  • Implements comprehensive error logging for operation monitoring
4.4 System Integration
The main application provides a unified interface through CLI commands with Clap [19] crate:
  • Server mode: quic-tun serve --dest <TCP SocketAddr> --bind <QUIC SocketAddr>
  • Client mode: quic-tun connect --dest <QUIC SocketAddr> --bind <TCP SocketAddr>
Notably, the implementation leverages Rust’s type system to ensure memory safety and prevent common concurrency issues. The use of asynchronous tasks throughout the implementation maximizes throughput while maintaining clear operational semantics.
All cryptographic operations default to Rustls’s [20] ring-based [21] provider, ensuring robust security primitives. The modular design facilitates straightforward adaptation to different tunneling requirements while maintaining consistent performance characteristics.
5 Experimental Methodology
This section presents a comprehensive experimental methodology to evaluate the performance characteristics of the TCP over QUIC tunneling implementation.
5.1 Experimental Setup
All experiments were conducted in a controlled virtual network environment to ensure repeatability and minimal external interference. The setup included the following components:
  • Virtual Machines: Two Debian 12.10.0 virtual machines deployed on the same physical host using KVM/QEMU virtualization. Each VM was configured with:
    • 4 vCPUs (with host-passthrough mode enabled for precise timing)
    • 8 GB RAM
  • Operating System and Kernel:
    • Debian GNU/Linux 12.10.0 (Bookworm)
    • Linux Kernel version 6.1 (default)
  • Unit Definitions:
    • Bandwidth: Measured in MB/s (megabytes per second)
[…]
  • Protocol Implementations:
    • TCP: Default Linux TCP stack using Cubic congestion control
    • QUIC: Implemented via the Quinn library (Rust)
  • Performance Measurement Tool: iperf3
    • iperf3 server was deployed on one VM and the client on the other
    • Tests were repeated 5 times for each protocol to obtain average values
    • Each test uses the default iperf3 parameter
    • The measured metrics included:
        Throughput (in Mbps)
        CPU utilization during transmission
  • Extra System Monitoring:
    • tcpdump was used to capture traffic for post-analysis and validation
5.2 Evaluation Metrics
The experimental framework measures:
  • Throughput:
    $\text{Throughput}=\frac{D}{t}$
    Where:
    • $D$ is the amount of data transferred (in bits)
    • $t$ is the transfer time (in seconds)
  • CPU Usage:
    CPU $\text{CPU Usage}=\frac{C\times P}{T}\times 100\%$
    where:
    • $C$: CPU cycles required to process one packet
    • $P$: Number of packets processed per second (packets/sec)
    • $T$: Total CPU cycles available per second (e.g., CPU frequency × number of cores)
When comparing tunneling technologies to direct connections, it is generally unnecessary to measure RTT (Round-Trip Time) in isolation. Any latency introduced by the tunnel will naturally manifest in end-to-end performance indicators such as throughput, application-level response time, or CPU utilization. Since RTT represents only one component of the overall performance footprint, separately testing it provides limited diagnostic value unless latency is a primary concern.
Furthermore, in scenarios where TCP traffic is encapsulated within a QUIC-based tunnel — despite being a non-standard configuration — RTT measurements taken at the TCP layer become methodologically unsound. This is because such RTT values are influenced by QUIC’s internal mechanisms, including congestion control, retransmission logic, and multiplexing of streams. Consequently, the observed RTT reflects a compound delay rather than the actual network-level round-trip time, making it unreliable for precise latency assessment or performance benchmarking.
6 Results & Analysis
6.1 Bandwidth Sensitivity
We firstly did a series on the effects of different bandwidths on the two connections, as shown in Table 1:
Table 1: Comparison of performance metrics under varying bandwidth
Bandwidth (MB/s) QUIC Tunnel Plain TCP
Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
100 132 126 0.3 17.6 824 820 6.5 15.3
10 76.0 70.1 0.5 5.8 76.5 73.2 0.6 1.2
1 11.8 6.18 0.3 1.1 11.1 7.75 0.3 0.2
It is not difficult to find that the throughput of TCP direct connection can completely crush the TCP over QUIC tunnel under excellent network conditions, but the TCP over QUIC tunnel itself is to improve the efficiency under weak network conditions, so in the following experiment, we use the bandwidth of 10MB/s by default.
6.2 Packet Loss Resilience
As illustrated in Table 2, We evaluated QUIC Tunnel and direct TCP at different packet loss rates using the default test parameters of iperf3. The packet loss rate is formally defined as:
Packet Loss $\text{Packet Loss Rate}=\frac{N_{\text{lost}}}{N_{\text{sent}}}\times 100\%$
where $N_{\text{lost}}$ denotes the number of lost packets, and $N_{\text{sent}}$ represents the total number of packets transmitted. This metric provides a standardized framework for assessing the system’s robustness in the presence of unreliable network conditions.
Table 2: Comparison of performance metrics under varying packet loss rates
Loss (%) QUIC Tunnel Plain TCP
Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
0 72.7 68.0 0.5 9.0 84.9 81.5 0.9 1.6
5 58.3 53.1 0.1 11.1 47.5 46.6 0.3 0.8
10 32.2 26.9 0.6 5.8 19.5 19.1 0.3 0.2
15 18.1 13.2 0.1 2.3 5.14 4.93 0.1 0.0
20 11.1 5.97 0.2 0.9 2.41 2.18 0.3 0.1
Figure 2: This figure shows the impact of the two connections on throughput at different packet loss rates
The results reveal several key observations:
  • Expected Early TCP Advantage: As expected, plain TCP slightly outperforms the QUIC tunnel at low packet loss rates (up to 5%) due to the overhead introduced by encapsulating traffic through dual protocol layers in the tunnel.
  • Consistent Throughput Advantage: But, Beyond 5% packet loss, the QUIC tunnel consistently outperforms direct TCP, maintaining throughput from 58.3 Mbps at 5% loss to 11.1 Mbps at 20% loss.
  • Graceful Performance Degradation: QUIC demonstrates more graceful degradation under increasing packet loss, whereas TCP performance degrades sharply (e.g., at 20% loss, QUIC achieves over 4.6$\times$ higher throughput).
  • Robustness through Design: These results highlight QUIC’s enhanced resilience, attributable to its native support for multiplexing, per-stream flow control, and more efficient loss recovery mechanisms.
6.3 Latency Impact Analysis
Table 3 presents the throughput degradation under increasing network latency, and the latency is usually defined as:
$\text{Latency}=\sum_{i=1}^{N}\left(T_{\text{prop}}^{i}+T_{\text{trans}}^{i}+T_{\text{queue}}^{i}+T_{\text{proc}}^{i}\right)$
Where:
  • $T_{\text{prop}}^{i}=\frac{d_{i}}{v_{i}}$: Propagation delay for hop $i$, where $d_{i}$ is the distance and $v_{i}$ is the propagation speed.
  • $T_{\text{trans}}^{i}=\frac{L_{i}}{R_{i}}$: Transmission delay for hop $i$, where $L_{i}$ is the packet length and $R_{i}$ is the transmission rate.
  • $T_{\text{queue}}^{i}$: Queuing delay for hop $i$, caused by waiting in the router’s queue, depending on network congestion and router buffer utilization.
  • $T_{\text{proc}}^{i}$: Processing delay for hop $i$, the time it takes for a router or switch to process a packet, including error checking, routing decision making, and forwarding.
In the context of our study, where we compare TCP over QUIC tunneling and direct TCP connections, we must account for the specific overheads introduced by the QUIC protocol. The latency for a TCP over QUIC tunnel, denoted as $\text{Latency}_{\text{TCP/QUIC}}$, can be described as follows:
$\text{Latency}=\sum_{i=1}^{N}\left(T_{\text{prop}}^{i}+T_{\text{trans}}^{i}+T_{\text{queue}}^{i}+T_{\text{proc}}^{i}\right)+T_{\text{QUIC}}^{i}$
Where $T_{\text{QUIC}}^{i}$ represents the additional latency introduced by the QUIC protocol itself at each hop. This latency includes another handshake overhead, encryption and decryption delays and multiplexing and stream management overhead.
Thus, the latency for TCP over QUIC tunneling is generally higher than direct TCP connections due to the added layers of security, multiplexing, and stream management, all of which contribute to the overall delay.
Table 3: Comparison of performance metrics under varying lags
Lag (ms) QUIC Tunnel Plain TCP
Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
0 77.6 70.3 0.8 8.1 66.1 53.3 0.6 0.6
50 59.7 54.7 0.3 6.8 61.8 56.1 1.2 0.8
100 59.4 53.8 0.3 5.2 50.7 37.8 0.6 0.9
200 37.0 31.5 0.3 3.4 43.2 37.2 0.7 0.6
500 15.5 9.97 0.1 1.5 4.72 3.23 0.0 0.0
1000 6.60 2.76 0.1 0.0 0.73 0.61 0.1 0.0
Figure 3: This figure shows the impact of the two connections on throughput at different lags
Observations:
  • Throughput Decrease with Latency: Both protocols exhibit the expected inverse relationship between latency and throughput. As latency increases, throughput decreases for both QUIC Tunnel and Plain TCP. While QUIC Tunnel maintains a higher throughput compared to Plain TCP at lower latencies (e.g., 0 ms and 50 ms), the performance gap narrows as latency increases. At higher latencies, both protocols experience significant throughput degradation.
  • Relative Stability of QUIC Tunnel: QUIC Tunnel demonstrates relatively more stable throughput at higher latencies. For instance, at 1000 ms, the throughput of QUIC Tunnel is still notably higher (6.60 Mbps on the sender side, 2.76 Mbps on the receiver side) compared to Plain TCP (0.73 Mbps and 0.61 Mbps, respectively). However, QUIC Tunnel’s performance is not immune to the impact of increasing latency.
  • CPU Usage Variability: CPU usage for both QUIC Tunnel and Plain TCP shows significant variability, likely due to factors such as system load and network conditions. While QUIC Tunnel tends to exhibit somewhat higher CPU usage at certain points (e.g., at 0 ms latency), it does not consistently outperform Plain TCP in terms of CPU usage. The observed differences in CPU usage between the two protocols are subject to random fluctuations and should not be used as a definitive measure of efficiency.
  • Impact of Latency on Plain TCP: Plain TCP shows a more pronounced decrease in throughput as latency increases, with a particularly sharp decline at latencies of 500 ms and beyond. In contrast, QUIC Tunnel’s throughput degradation is less severe under similar conditions, indicating that QUIC’s protocol design offers a degree of resilience to latency.
6.4 Packet Reordering Tolerance
The out-of-order rate is formally defined as:
[…]
where $N_{\text{out-of-order}}$ represents the number of packets received out of order, and $N_{\text{total}}$ is the total number of packets received. This metric provides insight into the system’s tolerance and recovery from packet reordering during transmission.
The out-of-order delivery tests from Table 4 reveal:
Table 4: Comparison of performance metrics under varying out of order rates
Out of Order (%) QUIC Tunnel Plain TCP
Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
Sender Receiver Sender Receiver Sender Receiver Sender Receiver
0 74.8 69.9 0.5 8.8 44.0 40.6 0.6 0.8
5 76.4 71.5 0.6 8.7 44.9 40.8 0.3 0.9
10 77.9 70.1 0.6 9.2 44.1 40.8 0.9 0.8
15 80.1 75.1 0.5 9.1 52.2 48.8 0.8 0.6
20 76.2 69.3 1.0 7.0 52.3 48.9 0.5 1.4
Figure 4: This figure shows the impact of the two connections on throughput at different out of order rates
And we can find that:
  • QUIC’s advantage in reordering: QUIC can handle packet reordering more effectively due to its stream-based, out-of-order tolerant design.
  • TCP over QUIC doesn’t fully inherit QUIC’s advantages: Despite being encapsulated in QUIC, TCP still relies on its in-order delivery and retransmission mechanisms, which can be affected by reordering.
  • Slightly improved stability over native TCP: TCP over QUIC may show slightly better stability compared to native TCP because QUIC can buffer and reorder packets more efficiently at the transport layer.
7 Conclusion & Future Work
This paper presents the implementation and evaluation of a TCP-based tunneling scheme over QUIC, with the aim of leveraging the advantages of QUIC while maintaining compatibility with traditional TCP-based applications. The results of comprehensive experiments demonstrate that under lossy or high-latency network conditions, TCP over QUIC offers significant resilience and performance advantages compared to native TCP connections. These benefits are primarily attributed to the inherent features of QUIC, including stream multiplexing and efficient loss recovery mechanisms.
However, our evaluation also reveals the limitations of TCP over QUIC in high-quality network environments. In such conditions, the dual-protocol stack introduces additional computational overhead and latency due to user-space processing, encryption, and protocol encapsulation. This overhead leads to lower throughput and higher CPU usage compared to direct TCP, which benefits from kernel-level optimizations. As such, while TCP over QUIC is highly effective in degraded or mobile network scenarios, it is less efficient in well-provisioned, stable networks.
Further research could explore more forms of QUIC-based network tunneling, such as TCP over WebTransport [22], to further develop the potential of QUIC in the transformation of the network protocol. In addition, TCP over QUIC tunnel can optimize the processing of different application layer protocols and improve the performance of traditional application layer protocols under this tunnel.

### v2

@@
 These findings provide insights into the trade-offs of TCP over QUIC tunneling and its suitability for deployment in dynamic or impaired networks.
-keywords:
-QUIC Protocol, TCP Encapsulation, Transport Layer Tunneling, Network Resilience
 1 Introduction
@@
 To address these challenges, the QUIC (Quick UDP Internet Connections) protocol, initially developed by Google and standardized in RFC 9000 [11], has emerged as a promising alternative.
-QUIC leverages UDP as its transport layer, enabling a 0-RTT handshake for known servers, integrating TLS 1.3 for enhanced security [5], and supporting stream multiplexing to mitigate HOL blocking.
+QUIC leverages UDP as its transport layer, enabling a 0-RTT handshake for known servers, integrating TLS 1.3 [18] for enhanced security [1], and supporting stream multiplexing to mitigate HOL blocking.
 These features make QUIC particularly suitable for modern web applications such as HTTP/3, where low latency and high security are critical.
@@
 In this work, we implement a simple TCP over QUIC tunneling system and evaluate its performance under diverse network conditions.
-Our contributions include a lightweight Rust-based implementation, a thorough comparison with native TCP, and a detailed analysis of throughput, latency, and robustness.
+We emphasize that this work is exploratory in nature, focusing on empirical evaluation rather than proposing a new tunneling mechanism.
+Our contributions include a lightweight Rust baseline implementation, a thorough comparison with native TCP, and a detailed analysis of throughput, latency, and robustness.
 The results provide new insights into the feasibility and advantages of using QUIC as a transport layer for TCP-based applications.
@@
 2.1 Comparison of QUIC and TCP Core Mechanisms
-The Transmission Control Protocol (TCP), as defined in RFC 9293 [1] and RFC 793 [2], has been fundamental for reliable Internet data transmission.
-However, QUIC, initially developed by Google and standardized in RFC 9000 [4], addresses TCP’s limitations, particularly for web applications requiring low latency and high security.
-A detailed comparison reveals:
-  • Connection Establishment: TCP employs a three-way handshake (SYN, SYN-ACK, ACK), requiring at least one RTT, with additional RTT for TLS 1.3 negotiation, as noted in Performance evaluation in TCP and QUIC [7].
-QUIC uses UDP as its transport layer, which enables a 0-RTT handshake for known servers.
-By merging the TCP and TLS handshakes into a single three-packet exchange, QUIC reduces latency and enhances initial data transmission speed.
-  • Multiplexing and Streams: TCP operates as a single ordered byte stream, susceptible to head-of-line (HOL) blocking, where packet loss delays all subsequent data.
-QUIC, as detailed in RFC 9000, supports arbitrary concurrent streams within a connection, each with independent flow control, mitigating HOL blocking and improving resource utilization [8].
-  • Security: TCP relies on separate TLS implementation post-handshake (RFC 8446), adding overhead.
-QUIC integrates TLS 1.3, encrypting all packets by default, reducing handshake latency and enhancing privacy, as discussed in QUIC vs TCP computational efficiency [9].
-  • Congestion Control: Both protocols implement congestion control, with TCP using algorithms like New Reno (RFC 6582) [10] and QUIC adopting similar mechanisms (RFC 9002) [11].
-However, QUIC’s user-space implementation allows for more agile updates, potentially offering better adaptability to network conditions.
-  • Transport Layer Implementation: TCP benefits from kernel-level optimizations, reusing state for efficiency.
-QUIC, running over UDP in user mode, sacrifices some kernel efficiency for flexibility, as highlighted in performance evaluations showing higher per-packet overhead due to user-kernel boundary crossings [9].
+TCP, defined in RFC 9293 [5] and RFC 793 [17], has long underpinned reliable Internet transport.
+QUIC, standardized in RFC 9000 [11], addresses TCP’s latency and security limitations, particularly for web applications.
+Key distinctions include:
+  • Connection Establishment: TCP requires a three-way handshake plus TLS 1.3 negotiation [15], incurring multiple RTTs. QUIC merges transport and TLS handshakes, enabling 0-RTT for known servers.
+  • Multiplexing: TCP’s single stream is prone to head-of-line blocking, while QUIC supports concurrent streams with independent flow control, reducing HOL delay [12].
+  • Security: TCP depends on separate TLS (RFC 8446), whereas QUIC integrates TLS 1.3, encrypting all packets and reducing latency [6].
+  • Congestion Control: Both employ algorithms such as New Reno (RFC 6582) [8] and QUIC-specific variants (RFC 9002) [10], with QUIC’s user-space design allowing faster evolution.
+  • Implementation: TCP benefits from kernel optimizations, while QUIC’s user-mode operation adds per-packet overhead but increases flexibility [6].
 2.2 Network Tunneling Principle
-Network tunneling, involves encapsulating one protocol within another to transport data across incompatible networks or enhance security [12], commonly used in virtual private networks (VPNs).
-For TCP over QUIC tunnels, the process entails wrapping TCP traffic within QUIC packets, leveraging QUIC’s features like reduced latency and integrated security.
-Two modes are relevant [13]:
-  • Datagram Mode: This mode, as proposed in Tunneling TCP inside QUIC [14], encapsulates entire packets, including network and transport headers, within QUIC datagrams.
-It supports any network-layer protocol but incurs significant overhead, with each packet requiring additional IP/UDP/QUIC headers, making it inefficient for TCP, especially with small packets.
-  • Stream Mode: Detailed in Tunneling TCP inside QUIC [14], this mode transports TCP bytestreams over QUIC streams, reducing overhead by eliminating redundant headers.
-It uses Type-Length-Value (TLV) messages for negotiation, such as TCP Connect TLV (0x00, 20 bytes) for destination specification, and supports flow control tuned to TCP’s receive window, enhancing efficiency for applications with frequent small packets.
-For this reason, we will adopt the stream mode in the implementation.
+Network tunneling encapsulates one protocol in another for compatibility or security [4]. For TCP over QUIC, two modes exist [16]:
+  • Datagram Mode: Encapsulates full packets (headers + payload) in QUIC datagrams [16], supporting arbitrary protocols but incurring high overhead.
+  • Stream Mode: Transmits TCP bytestreams over QUIC streams [16], reducing redundant headers via TLV negotiation (e.g., TCP Connect TLV) and aligning flow control with TCP windows.
+This mode is adopted in our implementation.
 3 Related Work
 3.1 QUIC and TCP Performance Studies
-The comparative performance of QUIC and TCP has been extensively studied in recent years, particularly in the context of web and mobile applications requiring low latency, high throughput, and strong reliability.
-QUIC, as a user-space transport protocol built atop UDP, introduces innovations such as multiplexed streams, integrated TLS 1.3, and 0-RTT connection establishment, positioning it as a successor to TCP for many latency-sensitive applications.
-Several studies have also focused on the efficiency of QUIC implementations in practice.
-They compared the CPU utilization of QUIC and TCP during high-load transmission, showing that QUIC consumes more computational resources due to user-space processing and encryption overhead.
-Nevertheless, this trade-off is often justified by QUIC’s resilience in lossy or fluctuating networks, where TCP performance degrades rapidly.
-In the context of tunneling, encapsulating TCP traffic over QUIC offers a hybrid solution that combines TCP’s widespread application compatibility with QUIC’s transport-layer robustness.
-This has motivated emerging research into TCP-over-QUIC systems that seek to leverage QUIC’s loss recovery, multiplexing, and security capabilities without altering the semantics of existing TCP applications.
-Our study extends this line of research by implementing and experimentally evaluating a practical TCP-over-QUIC tunnel using the Quinn library in Rust, further validating the protocol’s strengths and trade-offs under controlled packet loss and latency conditions.
-3.2 Comparison of different QUIC implementations
-Comparison of different QUIC implementations reveals notable differences in performance and characteristics.
-According to the work by Salim Hertelli et al.
-[15], various QUIC implementations such as Chromium-quic ”Quiche”, Cloudflare ”Quiche”, Mvfst, Msquic, Quic-go, Aioquic, and Haskell-quic have been developed in diverse programming languages like C/C++, Rust, Go, Python, and Haskell.
-These implementations support different versions of QUIC and employ various congestion control algorithms, including CUBIC, NewReno, and HyStart++.
-While all adhere to IETF specifications, they exhibit significant variations in project status, maintenance, supported features, and performance metrics.
-In a more detailed benchmarking study by Kevin Marti [16], MsQuic and Quinn were compared under different networking conditions.
-The results indicated that increasing the segment size for Quinn can significantly enhance its performance, up to doubling it in some cases.
-This suggests that Quinn benefits from larger segment sizes, which can reduce the overhead associated with frequent system calls and data copying.
-In fact, the patched version of Quinn, which sends larger segments, demonstrated improved throughput and fewer system calls compared to its unpatched version.
-This highlights a potential advantage of Quinn in scenarios where segment size can be optimized.
-However, when packet loss or reordering was introduced, MsQuic consistently outperformed Quinn by at least 50%.
-This indicates that MsQuic has more robust mechanisms for handling network instability.
-Despite this, Quinn’s ability to achieve high performance with optimized segment sizes and reduced system calls suggests that it can be highly efficient under stable network conditions.
-Overall, while MsQuic demonstrates superior resilience to network instability, Quinn’s performance can be significantly enhanced through specific optimizations, highlighting the importance of implementation-specific tuning in QUIC deployments.
-Therefore, we will use Quinn’s QUIC implementation to complete the TCP over QUIC tunnel construction.
+The performance of QUIC and TCP has been widely studied, especially for web and mobile applications requiring low latency and high throughput.
+QUIC, built atop UDP, integrates TLS 1.3, multiplexing, and 0-RTT setup, making it a strong TCP successor for latency-sensitive scenarios.
+Studies show QUIC incurs higher CPU overhead due to user-space processing and encryption, yet performs better than TCP in lossy or variable networks.
+For tunneling, encapsulating TCP over QUIC combines TCP’s compatibility with QUIC’s robustness.
+Recent work explores such hybrid systems to exploit QUIC’s loss recovery and multiplexing while preserving TCP semantics.
+Our study extends this line by implementing and experimentally evaluating a TCP over QUIC tunnel in Rust (Quinn), validating performance trade-offs under controlled loss and latency.
+3.2 Comparison of QUIC Implementations
+QUIC implementations vary in language, features, and performance.
+Hertelli et al.
+[9] compare Chromium-quic “Quiche”, Cloudflare “Quiche”, Mvfst, MsQuic, Quic-go, Aioquic, and Haskell-quic, noting differences in maintenance, congestion control (e.g., CUBIC, NewReno, HyStart++), and adherence to IETF standards.
+Marti [13] benchmarked MsQuic and Quinn.
+Optimizing Quinn with larger segment sizes nearly doubled throughput and reduced system calls, but MsQuic consistently outperformed Quinn under loss or reordering.
+These results highlight Quinn’s efficiency under stable networks and MsQuic’s robustness under instability.
+Given Quinn’s strengths and its Rust ecosystem integration, we adopt it for our TCP over QUIC tunnel.
 3.3 Identified Deficiencies
-Currently, there is no mature or standardized implementation of TCP over QUIC written in Rust.
-While some experimental projects exist, they are often incomplete or tailored to narrow use cases.
-In this work, rather than aiming for a production-ready or fully optimized solution, we present a minimal and functional implementation intended primarily for empirical evaluation.
-The goal is to provide a foundation for systematic performance testing and to better understand the behavior and trade-offs of TCP over QUIC tunneling under various network conditions.
-Moreover, there is a noticeable lack of comprehensive and methodologically sound performance evaluations for TCP over QUIC tunnel architectures.
-Prior studies have largely focused on comparing QUIC and TCP as standalone transport protocols, without exploring the unique characteristics and trade-offs introduced by encapsulating TCP over QUIC.
-In particular, metrics such as CPU overhead, throughput under varying packet loss rates, and performance in high-latency environments have not been systematically addressed in the context of tunneling.
-This paper aims to fill these gaps by implementing a practical TCP over QUIC tunnel in Rust and conducting targeted performance evaluations across a range of network conditions.
+Although there have been some related studies such as rstun [14], their design mainly emphasizes feature richness and product usability.
+However, for the purposes of controlled experiments, this relative complexity introduces additional variables that may obscure the essential performance characteristics of TCP over QUIC tunnels.
+Therefore, we developed a minimal but fully functional Rust-based baseline approach stripped of unnecessary mechanisms to ensure clarity and repeatability.
+On top of this baseline, we conducted systematic evaluations across diverse network conditions.
+Unlike prior work, which primarily compared QUIC and TCP directly, our study isolates the tunneling layer itself, allowing us to quantify trade-offs such as CPU overhead, throughput under loss, and robustness in high-latency environments.
 4 System Design & Implementation
-Figure 1: Improved Principle of QUIC-Tun: Showing the flow of network traffic from a local TCP client through a QUIC tunnel to a remote TCP server.
-This section delineates the implementation details of the TCP over QUIC tunneling. The Figure 1 shows the data flow through the tunnel. The implementation comprises three core components:
-  • A data transfer utility facilitating bidirectional communication
-  • A server-side QUIC tunnel endpoint
-  • A client-side TCP to QUIC proxy
-The system was implemented in Rust leveraging the Tokio [17] asynchronous runtime and Quinn [18] QUIC library.
-4.1 Bidirectional Data Transfer Mechanism
-The core data transfer functionality is encapsulated in the bidirectional_copy utility function.
-This asynchronous function establishes a full-duplex communication channel between a TCP stream and a QUIC connection by:
-  1. 1.
-    Splitting the TCP stream into separate reader and writer using tokio::net::TcpStream::split()
-  2. 2.
-    Creating two parallel asynchronous tasks
-    • Continuously copies data from TCP to QUIC send stream until EOF
-    • Relays incoming QUIC data to the TCP writer
-  3. 3.
-    Employing tokio::try_join! macro to synchronize both tasks while propagating any errors
-  4. 4.
-    Implementing proper stream shutdown procedures upon completion or failure
-The implementation ensures atomic operation through Rust’s ownership system and automatic resource management.
-4.2 Server-Side
-The QUIC server implementation operates as a persistent tunnel endpoint with the following architectural characteristics:
-4.2.1 Connection Handling
-  • Utilizes Quinn’s Endpoint abstraction to listen for incoming QUIC connections
-  • Spawns independent Tokio tasks for each connection to ensure isolation
-  • Implements continuous stream acceptance through accept_bi() for bidirectional communication
-4.2.2 Configuration Parameters
-  • Generates self-signed certificates for development purposes via rcgen.
-  • Configures transport parameters including:
-    • Maximum 100 concurrent bidirectional streams
-    • Disabled unidirectional streams
-    • 2-second keep-alive interval
-  • Maintains compatibility with standard TLS certificate verification.
-4.2.3 Tunnel Establishment
-  • For each accepted QUIC stream, establishes a new TCP connection to the target service
-  • Delegate data transfer to the bidirectional_copy utility
-  • Implements robust error handling for connection timeouts and application-level closures
-4.3 Client-Side
-The client component functions as a TCP to QUIC proxy with the following implementation details:
-4.3.1 Connection Initiation
-  • Binds to specified local TCP port using tokio::net::TcpListener
-  • Establish QUIC connection to remote endpoint up-on TCP client connection
-  • Opens new bidirectional QUIC stream for each TCP connection
-4.3.2 Security Configuration
-  • Supports configurable certificate verification:
-    • Standard TLS certificate chain validation
-    • Insecure mode with disabled server verification (for testing)
-  • Maintains identical keep-alive parameters (2-second interval) with server
-4.3.3 Data Forwarding
-  • Instantiates separate Tokio task per connection
-  • Utilizes same bidirectional_copy mechanism for consistent data transfer behavior
-  • Implements comprehensive error logging for operation monitoring
-4.4 System Integration
-The main application provides a unified interface through CLI commands with Clap [19] crate:
-  • Server mode: quic-tun serve --dest <TCP SocketAddr> --bind <QUIC SocketAddr>
-  • Client mode: quic-tun connect --dest <QUIC SocketAddr> --bind <TCP SocketAddr>
-Notably, the implementation leverages Rust’s type system to ensure memory safety and prevent common concurrency issues.
-The use of asynchronous tasks throughout the implementation maximizes throughput while maintaining clear operational semantics.
-All cryptographic operations default to Rustls’s [20] ring-based [21] provider, ensuring robust security primitives.
-The modular design facilitates straightforward adaptation to different tunneling requirements while maintaining consistent performance characteristics.
+Figure 1: QUIC-Tun architecture: traffic flow from a local TCP client through the QUIC tunnel to a remote TCP server.
+The TCP over QUIC tunneling system consists of three core components: a bidirectional data transfer utility, a server-side tunnel endpoint, and a client-side TCP–QUIC proxy.
+The implementation is in Rust using the Tokio runtime [21] and the Quinn QUIC library [20].
+4.1 Bidirectional Transfer
+The central bidirectional_copy utility establishes full-duplex forwarding between TCP and QUIC streams.
+It runs two concurrent tasks to relay traffic in both directions, synchronizes their lifecycle, and ensures clean shutdown.
+Rust’s ownership model and async runtime guarantee atomicity and safe resource management.
+4.2 Server-Side Endpoint
+The QUIC server operates as a persistent endpoint.
+Using Quinn’s Endpoint, it accepts connections and spawns isolated async tasks.
+Each QUIC stream is mapped to a TCP connection via the transfer utility.
+Configuration includes self-signed certificates (via rcgen), a maximum of 100 concurrent bidirectional streams, disabled unidirectional streams, and a 2s keep-alive interval.
+TLS compatibility is preserved.
+4.3 Client-Side Proxy
+The client listens on a local TCP port and establishes a QUIC connection upon demand.
+Each TCP stream is mapped to a bidirectional QUIC stream.
+Security options include standard TLS verification or a relaxed mode (for testing).
+Keep-alive parameters match the server.
+Each session runs in an async task with error logging.
+4.4 Integration
+A CLI interface (via Clap [19]) provides unified commands:
+  • Server: quic-tun serve --dest <TCP_ADDR> --bind <QUIC_ADDR>
+  • Client: quic-tun connect --dest <QUIC_ADDR> --bind <TCP_ADDR>
+Cryptographic operations rely on Rustls [22] with the ring provider [3]. The modular design supports extensibility to other tunneling scenarios while maintaining throughput and reliability.
 5 Experimental Methodology
-This section presents a comprehensive experimental methodology to evaluate the performance characteristics of the TCP over QUIC tunneling implementation.
+This section outlines the methodology used to evaluate TCP over QUIC tunneling.
 5.1 Experimental Setup
-All experiments were conducted in a controlled virtual network environment to ensure repeatability and minimal external interference. The setup included the following components:
-  • Virtual Machines: Two Debian 12.10.0 virtual machines deployed on the same physical host using KVM/QEMU virtualization. Each VM was configured with:
-    • 4 vCPUs (with host-passthrough mode enabled for precise timing)
-    • 8 GB RAM
-  • Operating System and Kernel:
-    • Debian GNU/Linux 12.10.0 (Bookworm)
-    • Linux Kernel version 6.1 (default)
-  • Unit Definitions:
-    • Bandwidth: Measured in MB/s (megabytes per second)
-[…]
+Experiments were conducted in a controlled virtual environment for repeatability. The setup included:
+  • Virtual Machines: Two Debian 12.10.0 VMs on the same host, each with 4 vCPUs (host-passthrough) and 8 GB RAM.
+  • OS/Kernel: Debian GNU/Linux 12.10.0 with Linux 6.1.
   • Protocol Implementations:
@@
     • QUIC: Implemented via the Quinn library (Rust)
-  • Performance Measurement Tool: iperf3
-    • iperf3 server was deployed on one VM and the client on the other
-    • Tests were repeated 5 times for each protocol to obtain average values
-    • Each test uses the default iperf3 parameter
-    • The measured metrics included:
-        Throughput (in Mbps)
-        CPU utilization during transmission
-  • Extra System Monitoring:
-    • tcpdump was used to capture traffic for post-analysis and validation
+  • Performance Tool: iperf3, repeated 5 times per protocol with default parameters, measuring throughput (Mbps) and CPU usage.
+  • Monitoring: tcpdump for traffic capture and validation.
+Units: Bandwidth in MB/s, throughput in Mbps ($1~\text{MB/s}=8~\text{Mbps}$).
 5.2 Evaluation Metrics
-The experimental framework measures:
-  • Throughput:
-    $\text{Throughput}=\frac{D}{t}$
-    Where:
-    • $D$ is the amount of data transferred (in bits)
-    • $t$ is the transfer time (in seconds)
-  • CPU Usage:
-    CPU $\text{CPU Usage}=\frac{C\times P}{T}\times 100\%$
-    where:
-    • $C$: CPU cycles required to process one packet
-    • $P$: Number of packets processed per second (packets/sec)
-    • $T$: Total CPU cycles available per second (e.g., CPU frequency × number of cores)
-When comparing tunneling technologies to direct connections, it is generally unnecessary to measure RTT (Round-Trip Time) in isolation.
-Any latency introduced by the tunnel will naturally manifest in end-to-end performance indicators such as throughput, application-level response time, or CPU utilization.
-Since RTT represents only one component of the overall performance footprint, separately testing it provides limited diagnostic value unless latency is a primary concern.
-Furthermore, in scenarios where TCP traffic is encapsulated within a QUIC-based tunnel — despite being a non-standard configuration — RTT measurements taken at the TCP layer become methodologically unsound.
-This is because such RTT values are influenced by QUIC’s internal mechanisms, including congestion control, retransmission logic, and multiplexing of streams.
-Consequently, the observed RTT reflects a compound delay rather than the actual network-level round-trip time, making it unreliable for precise latency assessment or performance benchmarking.
+  • Throughput: $D/t$, where $D$ is data (bits), $t$ is transfer time (s).
+  • CPU Usage: $(C\times P)/T\times 100\%$, where $C$ = CPU cycles per packet, $P$ = packets/sec, $T$ = total CPU cycles/sec.
+Measuring RTT in this context is unnecessary because tunnel effects appear in throughput, response time, or CPU usage.
+Moreover, TCP-layer RTT within a QUIC tunnel is distorted by QUIC’s congestion control and retransmission, making it unsuitable as a standalone metric.
 6 Results & Analysis
 6.1 Bandwidth Sensitivity
-We firstly did a series on the effects of different bandwidths on the two connections, as shown in Table 1:
 Table 1: Comparison of performance metrics under varying bandwidth
@@
 Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
-Sender Receiver Sender Receiver Sender Receiver Sender Receiver
+Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
 100 132 126 0.3 17.6 824 820 6.5 15.3
@@
 1 11.8 6.18 0.3 1.1 11.1 7.75 0.3 0.2
-It is not difficult to find that the throughput of TCP direct connection can completely crush the TCP over QUIC tunnel under excellent network conditions, but the TCP over QUIC tunnel itself is to improve the efficiency under weak network conditions, so in the following experiment, we use the bandwidth of 10MB/s by default.
+Table 1 compares throughput under varying bandwidth.
+As expected, native TCP outperforms the tunnel under ideal conditions, benefiting from kernel-level optimizations.
+Since the tunnel is designed for degraded networks, subsequent experiments adopt a default bandwidth of 10 MB/s.
 6.2 Packet Loss Resilience
-As illustrated in Table 2, We evaluated QUIC Tunnel and direct TCP at different packet loss rates using the default test parameters of iperf3. The packet loss rate is formally defined as:
-Packet Loss $\text{Packet Loss Rate}=\frac{N_{\text{lost}}}{N_{\text{sent}}}\times 100\%$
-where $N_{\text{lost}}$ denotes the number of lost packets, and $N_{\text{sent}}$ represents the total number of packets transmitted.
-This metric provides a standardized framework for assessing the system’s robustness in the presence of unreliable network conditions.
+We evaluated both approaches at different packet loss rates using iperf3. The packet loss rate is defined as:
+$\text{Packet Loss Rate}=\frac{N_{\text{lost}}}{N_{\text{sent}}}\times 100\%.$
 Table 2: Comparison of performance metrics under varying packet loss rates
@@
 Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
-Sender Receiver Sender Receiver Sender Receiver Sender Receiver
+Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
 0 72.7 68.0 0.5 9.0 84.9 81.5 0.9 1.6
@@
 Figure 2: This figure shows the impact of the two connections on throughput at different packet loss rates
-The results reveal several key observations:
-  • Expected Early TCP Advantage: As expected, plain TCP slightly outperforms the QUIC tunnel at low packet loss rates (up to 5%) due to the overhead introduced by encapsulating traffic through dual protocol layers in the tunnel.
-  • Consistent Throughput Advantage: But, Beyond 5% packet loss, the QUIC tunnel consistently outperforms direct TCP, maintaining throughput from 58.3 Mbps at 5% loss to 11.1 Mbps at 20% loss.
-  • Graceful Performance Degradation: QUIC demonstrates more graceful degradation under increasing packet loss, whereas TCP performance degrades sharply (e.g., at 20% loss, QUIC achieves over 4.6$\times$ higher throughput).
-  • Robustness through Design: These results highlight QUIC’s enhanced resilience, attributable to its native support for multiplexing, per-stream flow control, and more efficient loss recovery mechanisms.
-6.3 Latency Impact Analysis
-Table 3 presents the throughput degradation under increasing network latency, and the latency is usually defined as:
-$\text{Latency}=\sum_{i=1}^{N}\left(T_{\text{prop}}^{i}+T_{\text{trans}}^{i}+T_{\text{queue}}^{i}+T_{\text{proc}}^{i}\right)$
-Where:
-  • $T_{\text{prop}}^{i}=\frac{d_{i}}{v_{i}}$: Propagation delay for hop $i$, where $d_{i}$ is the distance and $v_{i}$ is the propagation speed.
-  • $T_{\text{trans}}^{i}=\frac{L_{i}}{R_{i}}$: Transmission delay for hop $i$, where $L_{i}$ is the packet length and $R_{i}$ is the transmission rate.
-  • $T_{\text{queue}}^{i}$: Queuing delay for hop $i$, caused by waiting in the router’s queue, depending on network congestion and router buffer utilization.
-  • $T_{\text{proc}}^{i}$: Processing delay for hop $i$, the time it takes for a router or switch to process a packet, including error checking, routing decision making, and forwarding.
-In the context of our study, where we compare TCP over QUIC tunneling and direct TCP connections, we must account for the specific overheads introduced by the QUIC protocol.
-The latency for a TCP over QUIC tunnel, denoted as $\text{Latency}_{\text{TCP/QUIC}}$, can be described as follows:
-$\text{Latency}=\sum_{i=1}^{N}\left(T_{\text{prop}}^{i}+T_{\text{trans}}^{i}+T_{\text{queue}}^{i}+T_{\text{proc}}^{i}\right)+T_{\text{QUIC}}^{i}$
-Where $T_{\text{QUIC}}^{i}$ represents the additional latency introduced by the QUIC protocol itself at each hop.
-This latency includes another handshake overhead, encryption and decryption delays and multiplexing and stream management overhead.
-Thus, the latency for TCP over QUIC tunneling is generally higher than direct TCP connections due to the added layers of security, multiplexing, and stream management, all of which contribute to the overall delay.
+Table 2 and Fig. 2 show that:
+  • TCP is slightly better at low loss ($\leq$5%) due to tunnel overhead.
+[…]
+  • QUIC degrades more gracefully, highlighting its efficient recovery mechanisms.
+The tunnel incurs somewhat higher CPU usage, mainly from encryption, user–kernel transitions, and encapsulation.
+6.3 Latency Impact
+Latency is defined as:
+$\text{Latency}=\sum_{i=1}^{N}(T_{\text{prop}}^{i}+T_{\text{trans}}^{i}+T_{\text{queue}}^{i}+T_{\text{proc}}^{i}),$
+with an extra $T_{\text{QUIC}}$ term for tunnel overhead.
 Table 3: Comparison of performance metrics under varying lags
@@
 Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
-Sender Receiver Sender Receiver Sender Receiver Sender Receiver
+Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
 0 77.6 70.3 0.8 8.1 66.1 53.3 0.6 0.6
@@
 Figure 3: This figure shows the impact of the two connections on throughput at different lags
-Observations:
-  • Throughput Decrease with Latency: Both protocols exhibit the expected inverse relationship between latency and throughput.
-As latency increases, throughput decreases for both QUIC Tunnel and Plain TCP.
-While QUIC Tunnel maintains a higher throughput compared to Plain TCP at lower latencies (e.g., 0 ms and 50 ms), the performance gap narrows as latency increases.
-At higher latencies, both protocols experience significant throughput degradation.
-  • Relative Stability of QUIC Tunnel: QUIC Tunnel demonstrates relatively more stable throughput at higher latencies.
-For instance, at 1000 ms, the throughput of QUIC Tunnel is still notably higher (6.60 Mbps on the sender side, 2.76 Mbps on the receiver side) compared to Plain TCP (0.73 Mbps and 0.61 Mbps, respectively).
-However, QUIC Tunnel’s performance is not immune to the impact of increasing latency.
-  • CPU Usage Variability: CPU usage for both QUIC Tunnel and Plain TCP shows significant variability, likely due to factors such as system load and network conditions.
-While QUIC Tunnel tends to exhibit somewhat higher CPU usage at certain points (e.g., at 0 ms latency), it does not consistently outperform Plain TCP in terms of CPU usage.
-The observed differences in CPU usage between the two protocols are subject to random fluctuations and should not be used as a definitive measure of efficiency.
-  • Impact of Latency on Plain TCP: Plain TCP shows a more pronounced decrease in throughput as latency increases, with a particularly sharp decline at latencies of 500 ms and beyond.
-In contrast, QUIC Tunnel’s throughput degradation is less severe under similar conditions, indicating that QUIC’s protocol design offers a degree of resilience to latency.
+From Table 3 and Fig. 3:
+  • Throughput decreases with latency for both protocols, but the tunnel maintains higher throughput at moderate latencies.
+  • At 1000 ms, the tunnel is still clearly ahead (6.60 vs. 0.73 Mbps on sender side).
+  • CPU usage varies, with the tunnel often slightly higher, but not consistently better or worse.
 6.4 Packet Reordering Tolerance
-The out-of-order rate is formally defined as:
-[…]
-where $N_{\text{out-of-order}}$ represents the number of packets received out of order, and $N_{\text{total}}$ is the total number of packets received.
-This metric provides insight into the system’s tolerance and recovery from packet reordering during transmission.
-The out-of-order delivery tests from Table 4 reveal:
+Out-of-order rate is defined as:
+$\text{Out-of-Order Rate}=\frac{N_{\text{out-of-order}}}{N_{\text{total}}}\times 100\%.$
 Table 4: Comparison of performance metrics under varying out of order rates
@@
 Bitrate (Mbps) CPU (%) Bitrate (Mbps) CPU (%)
-Sender Receiver Sender Receiver Sender Receiver Sender Receiver
+Sndr Rcvr Sndr Rcvr Sndr Rcvr Sndr Rcvr
 0 74.8 69.9 0.5 8.8 44.0 40.6 0.6 0.8
@@
 Figure 4: This figure shows the impact of the two connections on throughput at different out of order rates
-And we can find that:
-  • QUIC’s advantage in reordering: QUIC can handle packet reordering more effectively due to its stream-based, out-of-order tolerant design.
-  • TCP over QUIC doesn’t fully inherit QUIC’s advantages: Despite being encapsulated in QUIC, TCP still relies on its in-order delivery and retransmission mechanisms, which can be affected by reordering.
-  • Slightly improved stability over native TCP: TCP over QUIC may show slightly better stability compared to native TCP because QUIC can buffer and reorder packets more efficiently at the transport layer.
+As shown in Table 4 and Fig. 4:
+  • QUIC tolerates reordering well due to its stream-based design.
+  • TCP over QUIC inherits some benefit but remains constrained by TCP’s in-order semantics.
+  • The tunnel shows slightly better stability than native TCP under reordering.
+Finally, note that both inner TCP and outer QUIC implement congestion control independently.
+Their interaction may cause inefficiencies such as oscillation, which require further investigation beyond our VM-based tests.
 7 Conclusion & Future Work
@@
 As such, while TCP over QUIC is highly effective in degraded or mobile network scenarios, it is less efficient in well-provisioned, stable networks.
+Our current evaluation is limited to a single QUIC implementation (Quinn) and VM-based setups.
+Broader validation across diverse QUIC stacks and real-world WAN environments will further strengthen the conclusions.
 Further research could explore more forms of QUIC-based network tunneling, such as TCP over WebTransport [7], to further develop the potential of QUIC in the transformation of the network protocol.
 In addition, TCP over QUIC tunnel can optimize the processing of different application layer protocols and improve the performance of traditional application layer protocols under this tunnel.
+Future work should also investigate the congestion control interactions between the inner TCP and the outer QUIC layers, as this remains a fundamental challenge for tunneling designs.
+8 Data Availability
+The data sets generated and analyzed during the current study are not persistently stored, as all experimental data were dynamically generated during runtime.
+Due to variations in network conditions, system load, and inherent randomness, storing raw data such as packet captures or instantaneous performance metrics was deemed unnecessary.
+To ensure reproducibility, all testing methods and the full source code used in this study are publicly available at the GitHub repository: https://github.com/ElaBosak233/quic-tun.
+Researchers can reproduce the results by building and executing the code under similar or customized network conditions.